Saturday, August 31, 2013

On the telescope - Phase 1

AdSec-DX went on the telescope in its hub as planned on Aug 21. It looked like a nice little telescope on top of a big one, as you can see below...

The structure mounted on the hub of the secondary mirror is supporting a retroreflector which was used for all the tests made over the last week to recalibrate the AdSec: it was for the Thin Shell #3 (TS3) the first time it would be checked and characterized as a mirror for an active secondary.

The tests and calibration were made with a 4D laser interferometer. The series of three pictures below shows the evolution from the initial pupil (on the left) to what was achieved with 100 modes (on the right).

It became cleat that the mouse hole was likely going to be a problem. The area around it requires too high forces to be flattened. A close-up on the central hole and its mouse hole is seen below. 


A first pass at 600 modes while masking a central area slightly larger than the central obscuration of the telescope produced a mirror flattened as shown below.

The feature at the top-left external edge is due to a sequence of three clustered not working (therefore disabled) actuators. The previously known circular defect (groove) close to the external edge is also evident. The groove is about  5mm thick (FWHM) and about 100nm deep (WF). 

More tests were made and a lot of data were taken to prepare the simulation and preparatory work needed to use TS3 for AO corrections. A refined flattening scheme was devised which ended up with a flattened mirror shown below. It has a wave front error of 29nm rms, matching what was found for TS4 (the shell TS3 is replacing) and TS5 (the shell on the left side of the telescope).

Ad-Sec-Dx can now be used as a non-adaptive secondary mirror for seeing limited operation. The retroreflector holder has been dismounted and the mirror will be used for the coming nights of engineering where the telescope and the instruments will be brought back to life and checked out after nearly two months of shutdown.

Meanwhile, our colleagues in Arcetri will work on the AO software to slave the actuators which will be behind the masked area. The next, and final, step will be the calibration of the unit in closed loop for AO corrections. Work on the telescope sould start on Sep 13.

It is anticipated that TS3, which is only a spare shell, will not work as well as TS4. However, it should still allow excellent AO observations on the right side of the telescope while we wait for ts4 to be refurbished (not an easy thing to do...).

Another excellent team work with participants from Arcetri (on Mt Graham or remotely from Italy) and LBTO!

Saturday, August 17, 2013

AdSec-DX ready to go on the telescope

The goal of the last phase of AdSec-DX lab work was to end up with a functional adaptive secondary unit integrated in its hub and ready to go on the telescope.

It all started on Monday (Aug 12). After removing the shell mini-skirt, the Opticote layer was removed with the unit in its zenith pointing orientation. The mini-skirt was then reinstalled.

Opticote protection layer removal
Retention ring installation

On Tuesday, the central cover was installed and the internal spacers adjusted to reach the goal of ~400 µm gap everywhere. The installation was done first in zenith pointing position, but then adjusted after having moved the unit in horizon pointing position to facilitate the measurement of the gap between cover and thin shell.

Central cover installation

The unit was sealed and readied for installation in the hub. So far, so good!  

A final check of the unit was performed, setting it flat and slewing the unit at different elevations. After a few successful slews from 90° to 20° elevation, the shell started to display signs of contamination between the shell and its reference in an area close to its outer edge, which would not go away after one more slew and a couple more attempts at setting the shell . 

The contamination is seen through the larger gap between the shell and the reference 
located at 8pm at the outer edge of the shell

A clear setback for all involved. Unabated, the team just moved on for two more days of work, back to checking the shell and the reference body for the source of contamination and reassembling it again with more testing to follow! 

What had just been installed the day before was removed, and the shell gently lowered on its foam tray, No obvious contaminant, like a 50µm size particulate, was found after a thorough examination of both the shell and the reference body. After a tedious removal of all visible FODs (like "foreign object debris"...), most of them around 10µm in size, the unit was reassembled and sealed again. 

By Thursday night, the team was successful in completing the integration of the AdSec and testing the thin shell's membrane integrity. The results from these tests were consistent with the measurements made during last month's AdSec testing campaign.

On Friday evening, after one more day of work, the unit was finally in its hub! This last step went smoothly, and here is the result...

The units in its hub, ready to go on its swing arm

The hub is scheduled to go back on the telescope on Wednesday, Aug 21.

Kudos to the AdSec team of the week: Guido Brusa, Juan-Carlos Guerra and Michael Lefebvre!